# Cost-aware Stopping for Bayesian Optimization

### v1

@@ 5 Conclusion
We develop the PBGI/LogEIPC stopping rule for Bayesian optimization with varying evaluation costs and adaptive stopping. Paired with either the PBGI or LogEIPC acquisition function, it (a) satisfies a theoretical guarantee bounding the expected cumulative cost (Theorem 1), and (b) shows strong empirical performance in terms of cost-adjusted simple regret (Section 4). We believe our framework can be further extended to settings involving noisy, multi-fidelity or batched evaluations, as well as alternative objective formulations—for instance, applying a sigmoid transformation to test error rather than a linear one, to reflect real-world considerations such as a user’s willingness to adopt models whose performance may change sharply once test error falls below a certain threshold.
Acknowledgments and Disclosure of Funding
QX thanks Nairen Cao for his helpful suggestions, and AT thanks James T. Wilson for helpful comments on stopping rules for Bayesian optimization. LC is funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. AT was supported by Cornell University, jointly via the Center for Data Science for Enterprise and Society, the College of Engineering, and the Ann S. Bowers College of Computing and Information Science. PF was supported by AFOSR FA9550-20-1-0351. ZS was supported by the NSF under grant no. CMMI-2307008.
Appendix A Theoretical Analysis and Calculations
@@
We treat the two policies—meaning, the two acquisition function + stopping rule pairs—together and write $\mathcal{F}_{t}=\sigma\!\left(\{x_{i},y_{i}\}_{i=1}^{t}\right)$ for the filtration generated by the observations. Since we are minimizing, the one–step improvement after iteration $t$ is $y_{1:t-1}^{*}-y_{1:t}^{*}$, where recall that $y_{1:t}=\min_{1\leq i\leq t}y_{i}$. Since we initialize at the prior mean, we have $\mathbb{E}[y_{1:1}^{*}]=\mu$ and the quantity $y_{1:1}^{*}-\min\nolimits_{x\in X}f(x)$ has finite expectation $U<\infty$.
Step 1: A point-wise lower bound on the expected improvement before stopping.
Denote the posterior expected improvement function as $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$. While stopping hasn’t occurred, meaning $t<\tau$, by the stopping criteria definition we have $\max_{x\in X}\alpha_{t}^{\mathrm{EI}}(x)/c(x)\geq 1$. Hence, there exists at least one point with $\alpha_{t}^{\mathrm{EI}}(x)/c(x)\geq 1$. We now argue for each policy.
@@
Thus for both policies, we have
@@
Step 2: The cumulative expected improvement is at most $U$.
Set $\Delta_{t}=y_{1:t-1}^{*}-y_{1:t}^{*}\geq 0$ for $t\leq\tau$. Conditional on $\mathcal{F}_{t-1}$ and the choice of $x_{t}$, we have
@@
Taking expectations and summing up to $\tau$, we get
$\mathbb{E}\left[\sum_{t=2}^{\tau}\alpha_{t}^{\mathrm{EI}}(x_{t})\right]=\mathbb{E}\left[\sum_{t=2}^{\tau}\Delta_{t}\right]=\mathbb{E}\left[y_{1}^{*}-y_{1:\tau}^{*}\right].$ (19)
@@
$\mathbb{E}\left[y_{1}^{*}-y_{1:\tau}^{*}\right]\leq\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right]=U.$ (20)
Step 3: Expected cumulative cost control.
Summing (17) over $t\leq\tau$, taking expectations, and applying the result of Step 2, we have
$\mathbb{E}\left[\sum_{t=2}^{\tau}c(x_{t})\right]\leq\mathbb{E}\left[\sum_{t=2}^{\tau}\alpha_{t}^{\mathrm{EI}}(x_{t})\right]\leq U.$ (21)
@@
$\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]=\mathbb{E}[c(x_{1})]+\mathbb{E}\left[\sum_{t=2}^{\tau}c(x_{t})\right]\leq C+U,$ (22)
@@
Extension to other acquisition functions.
From the proof, we see that this bound extends to any acquisition function that satisfies the condition (17). That is, the result holds more generally to acquisition functions beyond PBGI and LogEIPC, as long as each selected point ensures sufficient expected improvement relative to its cost before stopping at time $\tau$.
Implications for budget-constrained settings and PBGI-D.
Suppose the user specifies a total evaluation cost budget $B$, but there is no direct unit conversion between evaluation cost and objective function value. As discussed in Section 3.2, Theorem 1 suggests that when using PBGI or LogEIPC, scaling costs by a factor of $\lambda=U/(B-C)$ ensures that the expected cumulative cost remains within budget $B$. If this choice of $\lambda$ proves conservative in practice—leading to lower-than-anticipated spending—it can be paired with the PBGI-D acquisition function variant introduced by Xie et al. [25], which starts with $\lambda=\lambda_{0}$ and reduces $\lambda$ by half each time stopping is triggered. Setting $\lambda_{0}=U/(B-C)$ provides a principled initial value that aligns the initial fixed-$\lambda$ phase with the budget in expectation and thus ensures that the adaptive decay is triggered in expectation.
Appendix B Experimental Setup
All experiments are implemented based on BoTorch [4]. Each Bayesian optimization procedure is initialized with $2(d+1)$ random samples, where $d$ is the dimension of the search domain. For Bayesian regret experiments, we follow the standard practice to generate the initial random samples using a quasirandom Sobol sequence. For empirical experiments, we randomly sample configuration IDs from a fixed pool of candidates—2,000 for LCBench and 32,768 for NATS-Bench. All computations are performed on CPU.
@@ Appendix B Experimental Setup
For all experiments, we follow the standard practice to apply Matérn kernels with smoothness $5/2$ and length scales learned from data via maximum marginal likelihood optimization, and standardize input variables to be in $[0,1]^{d}$. For empirical experiments, we standardize output variables to be zero-mean and unit-variance, but not for Bayesian regret experiments. In this work, we consider the noiseless setting and set the fixed noise level to be $10^{-6}$. In the unknown-cost experiments, we follow Astudillo et al. [2] to model the objective and the logarithm of the cost function using independent Gaussian processes.
@@
For PBGI, we follow Xie et al. [25] to compute the Gittins indices using $100$ iterations of bisection search without any early stopping or other performance and reliability optimizations.
For UCB/LCB-based acquisition functions and stopping rules, we follow the original GP-UCB paper Srinivas et al. [22] and the choice in UCB/LCB based stopping rules [17, 11] to use the schedule $\beta_{t}=2\log(dt^{2}\pi^{2}/6\delta)$, where $d$ is the dimension. We also adopt their choice of $\delta=10^{-1}$ and a scale-down factor of $5$.
@@
For PRB, we follow Wilson [24] to use the schedule $N_{t}=\max(\lceil 64*1.5^{t-1}\rceil,1000)$ for number of posterior samples, risk tolerance $\delta=0.05$. The error bound $\epsilon$ is set to be 0.1 for Bayesian regret experiments and 0.5% of the best test error (here, the misclassification rate) among all configurations for the empirical experiments.
For the 8D Bayesian regret experiments, we apply moving average over 20 iterations to mitigate oscillations in the optimizer. Figure 6 illustrates the challenges these oscillations pose when computing stopping rule statistics and shows the improvement with moving average. For consistency, we also apply 20-iteration averaging to the GSS and Convergence baselines.
@@
We omit the PRB baseline for the 8D Bayesian regret experiment due to its high computational cost—requiring up to global optimum computations of 1000 sample paths per iteration. We also omit the KG acquisition function and stopping rule for the same reason, as they are shown to be computationally intensive in the runtime experiments of Xie et al. [25].
@@
[…]
In the periodic cost setting, the evaluation cost fluctuates across the domain. Following [2], we define the periodic cost as
[…]
@@
A visualization of the three cost functions is provided in Figure 7.
Figure 7: Surface plots of cost functions over $[0,1]^{2}$: (Left) Uniform cost of $1$ across domain. (Middle) Normalized linear cost increasing with the mean of $x_{1}$ and $x_{2}$. (Right) Periodic cost with $\alpha=2$, $\beta=2$, normalized by Bessel-based factor.
@@
Figure 8: Empirical relationship between the number of model parameters and runtime for three LCBench datasets. Each subplot shows a scatter plot of actual runtime ($y$-axis) against number of model parameters ($x$-axis), along with a fitted linear regression line. The observed linear trend supports using 0.001 times the number of model parameters as a proxy for runtime. For Fashion-MNIST and volkert, the fitted slopes are close to 0.001. The slope for higgs is slightly higher, possibly due to a few outliers.
@@
    • Number of Output Features: 1 (binary classification: income $>50K$ or $\leq 50K$)
@@
  • Volkert:
    • Input Dimension: 147 (the dataset consists of 147 features)
    • Number of Output Features: 10 (corresponding to 10 classes)
@@
[…]
@@
[…]
@@
Figure 9: Empirical relationship between the number of FLOPs and runtime for three NATS-Bench datasets. Each subplot shows a heatmap of actual runtime ($y$-axis) against number of FLOPs ($x$-axis), along with a fitted linear regression line. The observed linear trend supports using $\alpha\times\#\text{FLOPs}+\beta$ as a proxy for runtime.
@@
Appendix C Additional Experiment Results
@@ C.1 Runtime Comparison
C.2 Order of Stopping and Posterior Updates
Following the discussions in Section 3, we always compute our proposed stopping rule with respect to the optimal acquisition function value of the next round—namely, the one which is obtained after posterior updates have been performed. One could alternatively consider checking the stopping criteria before posterior updates, which is backward-looking rather than forward-looking. Figure 11 provides an empirical comparison between the two choices, showing that stopping after the posterior update leads to stronger empirical performance.
This suggests that the theoretical guarantee for the Gittins index policy in the correlated Pandora’s Box setting by Gergatsouli and Tzamos [10], which is based on the before-posterior-update stopping, could potentially be improved by adopting the after-posterior-update stopping.
Figure 11: Illustration of a single draw from a Matérn-$5/2$ Gaussian process on $[0,1]$ with lengthscale $0.1$, optimized using PBGI acquisition function under uniform cost and cost-scaling factor $\lambda=0.01$. We compare two variants of PBGI stopping rules: the before-posterior-update (this-round) stopping rule and the after-posterior-update (next-round) stopping rule. Left: The latent objective function (solid gray) and evaluation sequences for before-posterior-update stopping (blue circles) and after-posterior-update stopping (orange crosses). The dotted blue line and the dashed orange line mark the best observed value under each respective rule. Right: Cost-adjusted regret for each stopping rule. In this example, after-posterior-update stopping achieves strictly lower cost-adjusted regret despite performing more evaluations.
C.3 Additional Experiment Results: Empirical
@@
We count the number of trials in which a stopping rule fails to trigger within our iteration cap of 200 and present the results in Table 1. From the table, we observe that on datasets from the NATS benchmark, regret-based and acquisition-based stopping rules—except for ours—often fail to stop early. On LCBench datasets, some regret-based stopping rules such as SRGap-med and UCB-LCB also frequently exceed the cap. In contrast, our stopping rule consistently stops early, which aligns with our theoretical result in Theorem 1.
@@
Fashion-MNIST 0 0 47 50 0 0 0
higgs 0 0 50 50 0 0 0
volkert 0 0 50 50 0 0 0
Cifar10 0 48 50 50 0 0 49
Cifar100 0 50 50 50 0 0 50
ImageNet 0 50 50 50 0 0 45
@@
As shown in the empirical setting in Figure 12-14, cost-adjusted regret generally decreases in the early iterations and then increases. The turning point is exactly the hindsight-optimal fixed-iteration stopping point, and our PBGI/LogEIPC stopping rule consistently performs close to this optimum, particularly when paired with the PBGI acquisition function.
@@
Figure 14: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations on the volkert dataset. The objective function is the validation error, and the evaluation cost is the proxy runtime, scaled by three different cost-scaling factors $\lambda=10^{-3},10^{-4},10^{-5}$. We can see that the PBGI/LogEIPC stopping rule consistently achieves the cost-adjusted regret close to hindsight optimal adaptive stopping and hindsight optimal fixed-iteration stopping when paired with the PBGI or TS acquisition function, particularly with PBGI.
@@
In the known-cost setting of hyperparameter tuning, a practical approach is to use a proxy for runtime as the evaluation cost during the Bayesian optimization procedure. In our case, we use the number of model parameters scaled by a constant factor, which can be known in advance and has been shown to correlate well with the actual runtime. However, for reporting performance, one may prefer to use the actual runtime to better reflect real-world cost. To assess the impact of this cost model mismatch, we compare the cost-adjusted simple regret obtained when evaluation costs are computed using either the proxy runtime or the actual runtime. As shown in Figure 15 and Figure 16, our PBGI/LogEIPC stopping rule remains close to the hindsight optimal even when there is a mismatch, although its ranking may shift slightly (e.g., from best to second-best on the higgs dataset).
@@
Figure 16: Comparison of cost-adjusted simple regret on NATS-Bench with $\lambda=10^{-4}$, using scaled proxy runtime vs. scaled actual runtime as evaluation cost. Results are nearly identical to the cost-model-match setting.

### v2

@@ 5 Conclusion
We develop the PBGI/LogEIPC stopping rule for Bayesian optimization with varying evaluation costs. Paired with either the PBGI or LogEIPC acquisition function, it (a) satisfies a theoretical guarantee bounding the expected cost-adjusted simple regret (Section˜3.1), and (b) shows strong empirical performance in terms of cost-adjusted simple regret (Section˜4). We believe our framework can be extended to settings involving noisy, multi-fidelity or batched evaluations, as well as alternative objective formulations—for instance, applying a sigmoid transformation to test error rather than a linear one, to reflect real-world user preferences that shift sharply once error falls below a threshold.
Acknowledgments
QX thanks Nairen Cao for helpful discussions on hindsight optimal, Theodore Brown for discussions on LCBench runtime proxy, and Raul Astudillo for discussions on unknown cost handling for LogEIPC. AT thanks James T. Wilson for helpful comments on stopping rules for Bayesian optimization. LC is funded by the European Union (ERC-2022-SYG-OCEAN-101071601). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. AT was supported by Cornell University, jointly via the Center for Data Science for Enterprise and Society, the College of Engineering, and the Ann S. Bowers College of Computing and Information Science. PF was supported by AFOSR FA9550-20-1-0351. ZS was supported by the NSF under grant no. CMMI-2307008.
Appendix A LLM Usage Disclosure
We used large language models (LLMs) to assist with writing and editing this paper (e.g., improving clarity and readability of drafts). LLMs were also used to help polish proofs, to facilitate experiment scripting, and to generate plotting code. All technical content, research ideas, and final implementations were developed, verified, and approved by the authors.
Appendix B Theoretical Analysis and Calculations
In the following lemma, we prove a point-wise lower bound on the expected improvement before stopping for our recommended pairing of PBGI/LogEIPC stopping rule paired with either PBGI or LogEIPC acquisition function. [^4] In fact, any acquisition function (e.g., expected imporvement-cost (EIC) (Hu et al., 2025)) that ensures sufficient expected improvement relative to cost before the stopping time $\tau$ can apply here.
@@ Appendix B Theoretical Analysis and Calculations
While stopping has not occurred, meaning $t<\tau$, by the stopping criteria definition we have $\max_{x\in X}\alpha_{t}^{\mathrm{EI}}(x)/c(x)\geq 1$. Hence, there exists at least one point with $\alpha_{t}^{\mathrm{EI}}(x)/c(x)\geq 1$. We now argue for each algorithm.
@@
Thus for both algorithms, we have
@@
We can now use Lemma˜1 to prove the following theorem, where we show that our PBGI/LogEIPC stopping rule (paired with the PBGI or LogEIPC acquisition function) also achieves cost-adjusted simple regret no worse than a naive baseline—stopping-immediately (Immediate). Notably, this guarantee may not hold for other acquisition–stopping rule pairings. Moreover, in the worst case, this is the best guarantee we can hope for—for instance, the evaluation costs can be uniformly high and no point is worth evaluating.
See 2
Proof.
We treat the two algorithms—meaning, the two acquisition function and stopping rule pairs—together and write $\mathcal{F}_{t}=\sigma\!\left(\{x_{i},y_{i}\}_{i=1}^{t}\right)$ for the filtration generated by the observations. Since we are minimizing, the one–step improvement after iteration $t$ is $y_{1:t-1}^{*}-y_{1:t}^{*}$, where recall that $y_{1:t}^{*}=\min_{1\leq i\leq t}y_{i}$. Since we initialize at the prior mean, we have $\mathbb{E}[y_{1:1}^{*}]=\mu$ and the quantity $y_{1:1}^{*}-\min\nolimits_{x\in X}f(x)$ has finite expectation $U<\infty$. Denote the posterior expected improvement function as $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$.
By Lemma˜1, $c(x_{t})\leq\alpha_{t}^{\mathrm{EI}}(x_{t})$ for all $t\leq\tau$. Set $\Delta_{t}=y_{1:t-1}^{*}-y_{1:t}^{*}\geq 0$ for $2\leq t\leq\tau$. Conditional on $\mathcal{F}_{t-1}$ and the choice of $x_{t}$, we have
@@
Taking expectations and summing up from 2 to $\tau$, by optional stopping theorem we get
$\mathbb{E}\left[\sum_{t=2}^{\tau}\alpha_{t}^{\mathrm{EI}}(x_{t})\right]=\mathbb{E}\left[\sum_{t=2}^{\tau}\Delta_{t}\right]=\mathbb{E}\left[y_{1:1}^{*}-y_{1:\tau}^{*}\right].$ (16)
@@
$\mathbb{E}\left[y_{1:1}^{*}-y_{1:\tau}^{*}\right]\leq\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right]=U.$ (17)
Summing (14) over $t\leq\tau$, taking expectations, and applying (17), we have
$\mathbb{E}\left[\sum_{t=2}^{\tau}c(x_{t})\right]\leq\mathbb{E}\left[\sum_{t=2}^{\tau}\alpha_{t}^{\mathrm{EI}}(x_{t})\right]\leq\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right].$ (18)
@@
$\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]=\mathbb{E}[c(x_{1})]+\mathbb{E}\left[\sum_{t=2}^{\tau}c(x_{t})\right]\leq C+\mu-\mathbb{E}\left[\min_{x\in X}f(x)\right]=C+U,$ (19)
@@
Note 4.
In this paper, we focus on the setting where $f$ is drawn from a Gaussian process, i.e., $f\sim\mathcal{GP}(\mu,K)$. In this case, the term $U$ can be further bounded above using classical results on the expected supremum/infimum of Gaussian processes; see, for example, Lifshits (2012, Theorem 10.1).
See 3
Proof.
Since the Bayesian optimization algorithm considers the post-scaling cost, by Theorem˜2, $\mathbb{E}[\sum_{t=1}^{\tau}\lambda\cdot c(x_{t})]\leq\lambda C+U$. Since $\lambda=U/(B-C)$, we have
$\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq\frac{\lambda C+U}{\lambda}=C+\frac{U}{\lambda}=B.$
Appendix C Experimental Setup
All experiments are implemented based on BoTorch (Balandat et al., 2020). Each Bayesian optimization procedure is initialized with $2(d+1)$ random samples, where $d$ is the dimension of the search domain. For Bayesian regret experiments, we follow the standard practice to generate the initial random samples using a quasirandom Sobol sequence. For empirical experiments, we randomly sample configuration IDs from a fixed pool of candidates—2,000 for LCBench and 32,768 for NATS-Bench. All computations are performed on CPU.
@@ Appendix C Experimental Setup
For all experiments, we follow the standard practice to apply Matérn kernels with smoothness $5/2$ and length scales learned from data via maximum marginal likelihood optimization, and standardize input variables to be in $0,1]^{d}$. For empirical experiments, we standardize output variables to be zero-mean and unit-variance, but not for Bayesian regret experiments. In this work, we consider the noiseless setting and set the fixed noise level to be $10^{-6}$. In the unknown-cost experiments, we follow Astudillo et al. (2021) to model the objective and the logarithm of the cost function using independent Gaussian processes.
@@
For PBGI, we follow Xie et al. (2024) to compute the Gittins indices using $100$ iterations of bisection search without any early stopping or other performance and reliability optimizations.
For UCB/LCB-based acquisition functions and stopping rules, we follow the original GP-UCB paper Srinivas et al. (2009) and the choice in UCB/LCB based stopping rules (Makarova et al., 2022; Ishibashi et al., 2023) to use the schedule $\beta_{t}=2\log(dt^{2}\pi^{2}/6\delta)$, where $d$ is the dimension. We also adopt their choice of $\delta=10^{-1}$ and a scale-down factor of $5$.
@@
For PRB, we follow Wilson (2024) to use the schedule $N_{t}=\max(\lceil 64*1.5^{t-1}\rceil,1000)$ for number of posterior samples, risk tolerance $\delta=0.05$. The error bound $\epsilon$ is set to be 0.1 for Bayesian regret experiments and 0.5% of the best test error (here, the misclassification rate) among all configurations for the empirical experiments.
For the 8D Bayesian regret experiments, we apply moving average over 20 iterations to mitigate oscillations in the optimizer. Figure˜5 illustrates the challenges these oscillations pose when computing stopping rule statistics and shows the improvement with moving average. For consistency, we also apply 20-iteration averaging to the GSS and Convergence baselines.
@@
We omit the KG acquisition function and stopping rule due to its high computational cost, as they are shown to be computationally intensive in the runtime experiments of Xie et al. (2024).
@@
$\text{linear\_cost}(x)=\frac{1+20\cdot\left(\frac{1}{d}\sum_{i=1}^{d}x_{i}\right)}{11}.$
In the periodic cost setting, the evaluation cost fluctuates across the domain. Following Astudillo et al. (2021), we define the periodic cost as
$\text{periodic\_cost}(x)=\frac{\exp\left(\frac{\alpha}{d}\sum_{i=1}^{d}\cos\left(2\pi\beta(x_{i}-x^{*}_{i})\right)\right)}{\left[I_{0}\left(\frac{\alpha}{d}\right)\right]^{d}},$
@@
A visualization of the three cost functions is provided in Figure˜6.
Figure 6: Surface plots of cost functions over $0,1]^{2}$: (Left) Uniform cost of 11 across domain. (Middle) Normalized linear cost increasing with the mean of $x_{1}$ and $x_{2}$. (Right) Periodic cost with $\alpha=2$, $\beta=2$, normalized by Bessel-based factor.
@@
Figure 7: Empirical relationship between the number of model parameters and runtime for three LCBench datasets. Each subplot shows a scatter plot of actual runtime ($y$-axis) against number of model parameters ($x$-axis), along with a fitted linear regression line. The observed linear trend supports using 0.001 times the number of model parameters as a proxy for runtime. For Fashion-MNIST and adult, the fitted slopes are close to 0.001. The slope for higgs is slightly higher, possibly due to a few outliers.
@@
    • Number of Output Features: 1 (binary classification: income >50K>50K or $\leq 50K$)
@@
$\text{layer\_params}_{i\to i+1}$ $=\text{layer}_{i}\cdot\text{layer}_{i+1}+\text{layer}_{i+1}\quad\text{(weights + biases)}$ (20)
@@
where $\text{layer}_{0}=\text{input\_dim},\quad\text{layer}_{L}=\text{output\_feat}$ (22)
@@
Figure 8: Empirical relationship between the number of FLOPs and runtime for the three NATS-Bench datasets. Each subplot shows a heatmap of actual runtime ($y$-axis) against number of FLOPs ($x$-axis), along with a fitted linear regression line. The observed linear trend supports using $\alpha\times\#\text{FLOPs}+\beta$ as a proxy for runtime.
@@
Appendix D Additional Experiment Results
@@ D.1 Runtime Comparison
D.2 Order of Stopping and Posterior Updates
Following the discussions in Section˜3, we always compute our proposed stopping rule with respect to the optimal acquisition function value of the next round—namely, the one which is obtained after posterior updates have been performed. One could alternatively consider checking the stopping criteria before posterior updates, which is backward-looking rather than forward-looking. Figure˜10 provides an empirical comparison between the two choices, showing that stopping after the posterior update leads to stronger empirical performance.
This suggests that the theoretical guarantee for the Gittins index policy in the correlated Pandora’s Box setting by Gergatsouli & Tzamos (2023), which is based on the before-posterior-update stopping, could potentially be improved by adopting the after-posterior-update stopping.
Figure 10: Illustration of a single draw from a Matérn-$5/2$ Gaussian process on $[0,1]$ with lengthscale $0.1$, optimized using PBGI acquisition function under uniform cost and cost-scaling factor $\lambda=0.01$. We compare two variants of PBGI stopping rules: the before-posterior-update (this-round) stopping rule and the after-posterior-update (next-round) stopping rule. Left: The latent objective function (solid gray) and evaluation sequences for before-posterior-update stopping (blue circles) and after-posterior-update stopping (orange crosses). The dotted blue line and the dashed orange line mark the best observed value under each respective rule. Right: Cost‐adjusted regret for each stopping rule. In this example, after-posterior-update stopping achieves strictly lower cost-adjusted regret despite performing more evaluations.
D.3 Additional Experiment Results: Empirical
@@
To isolate the effect of the acquisition function on cost-adjusted regret, we report the simple regret of several acquisition functions in the budgeted setting. We compare LogEIPC, PBGI, LCB, and TS, and additionally include PBGI-D with our recommended choice $\lambda_{0}=U/(B-C)$ from Section˜3.1.1, where $U=50$ (reflecting the [0,100] range of classification accuracy) and BC=10, [^0] B-C=10{,}000 (corresponding to a budget after initial evaluation of 10,000 seconds, or roughly 3 hours). Cost-aware methods (LogEIPC and the PBGI variants) outperform cost-unaware ones (UCB and TS), with PBGI at smaller $\lambda$ consistently better than LogEIPC. This mirrors the findings of Xie et al. (2024) and helps explain the strong performance of the matched PBGI combination in our main experiments.
Figure 11: Comparison of simple regret of five acquisition functions: LogEIPC, PBGI($\lambda=10^{-4}$), LCB, TS, and PBGI-D on LCBench datasets, using scaled proxy runtime as evaluation cost. We can see that indeed the cost-aware LogEIPC and PBGI outperforms the cost-unaware UCB and TS. PBGI-D with out recommended $\lambda_{0}$ is also competitive.
@@
We count the number of trials in which a stopping rule fails to trigger within our iteration cap of 200 and present the results in Table˜1. From the table, we observe that on datasets from the NATS benchmark, regret-based and acquisition-based stopping rules—except for ours—often fail to stop early. On LCBench datasets, some regret-based stopping rules such as SRGap-med and UCB-LCB also frequently exceed the cap. In contrast, our stopping rule consistently stops early, which aligns with our theoretical result in Corollary˜3.
@@
Fashion-MNIST 0 0 0 50 0 0 0
adult 5 0 7 38 0 0 6
higgs 0 0 31 50 0 0 0
Cifar10 0 32 50 50 0 0 26
Cifar100 3 17 50 50 0 0 2
ImageNet 0 23 50 50 0 0 0
@@
As shown in the empirical setting in Figures˜12, 13, 14, 15, 16 and 17, cost-adjusted regret generally decreases in the early iterations and then increases. The turning point is exactly the hindsight-optimal fixed-iteration stopping point, and our PBGI/LogEIPC stopping rule consistently performs close to this optimum, particularly when paired with the PBGI acquisition function.
@@
Figure 13: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations on the adult dataset. The objective function is the validation error, and the evaluation cost is the proxy runtime, scaled by three different cost-scaling factors $\lambda=10^{-3},10^{-4},10^{-5}$. We can see that the PBGI/LogEIPC stopping rule consistently achieves the cost-adjusted regret close to hindsight optimal adaptive stopping and hindsight optimal fixed-iteration stopping when paired with the PBGI or TS acquisition function, particularly with PBGI.
@@
Figure 15: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations on the cifar10-valid dataset. The objective function is the validation error, and the evaluation cost is the proxy runtime, scaled by three different cost-scaling factors $\lambda=10^{-4},10^{-5},10^{-6}$. We can see that the PBGI/LogEIPC stopping rule consistently achieves the best cost-adjusted regret when paired with the LogEIPC, PBGI, or TS acquisition function, particularly with PBGI. These combinations not only approach the hindsight optimal adaptive stopping but also perform comparably to the hindsight optimal fixed-iteration stopping.
Figure 16: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations on the cifar100 dataset. The objective function is the validation error, and the evaluation cost is the proxy runtime, scaled by three different cost-scaling factors $\lambda=10^{-4},10^{-5},10^{-6}$. The PBGI/LogEIPC stopping rule remains competitive at $\lambda=10^{-4}$ and $10^{-6}$, though not always the best. At $\lambda=10^{-5}$, unlike in most experiments, it stops noticeably late (though still outperforming several other rules), even when paired with its matching acquisition function. By Lemma˜1, under model match, the PBGI/LogEIPC rule with the corresponding acquisition function should never incur negative expected cost-adjusted regret before stopping. The suboptimal behavior observed here points to significant model mismatch on the cifar100 dataset.
Figure 17: Comparison of cost-adjusted simple regret across acquisition function and stopping rule combinations on the ImageNet16-120 dataset. The objective function is the validation error, and the evaluation cost is the proxy runtime, scaled by three different cost-scaling factors $\lambda=10^{-4},10^{-5},10^{-6}$. The PBGI/LogEIPC stopping rule remains competitive at $\lambda=10^{-4}$ and $10^{-6}$, though not always the best. At $\lambda=10^{-5}$, unlike in most experiments, it stops noticeably late (though still outperforming several other rules), even when paired with its matching acquisition function. By Lemma˜1, under model match, the PBGI/LogEIPC rule with the corresponding acquisition function should never incur negative expected cost-adjusted regret before stopping. The suboptimal behavior observed here points to significant model mismatch on the ImageNet16-120 dataset.
@@
In the known-cost setting of hyperparameter tuning, a practical approach is to use a proxy for runtime as the evaluation cost during the Bayesian optimization procedure. In our case, we use the number of model parameters scaled by a constant factor, which can be known in advance and has been shown to correlate well with the actual runtime. However, for reporting performance, one may prefer to use the actual runtime to better reflect real-world cost. To assess the impact of this cost model mismatch, we compare the cost-adjusted simple regret obtained when evaluation costs are computed using either the proxy runtime or the actual runtime. As shown in Figure 18 and Figure LABEL:fig:cost_mismatch_nats, our PBGI/LogEIPC stopping rule remains close to the hindsight optimal even when there is a mismatch, although its ranking may shift slightly (e.g., from best to second-best on the higgs dataset).